OpenStudy (anonymous):

Mark and Jill each have a $6500 loan with a loan term of 4 years. The interest rate on Mark’s loan is 3% and the interest rate on Jill’s loan is 4%. Mark’s monthly payment is $144 and Jill’s monthly payment is $147. How much more in interest will Jill pay than Mark? A. $144 B. $412 C. $556 D. $968

OpenStudy (amistre64):

how do we determine the total cost of a loan?

OpenStudy (anonymous):

I dont know

OpenStudy (anonymous):

i need help with this

OpenStudy (amistre64):

well, if we are paying for something, then the amount we pay is how much it costs right?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

then 4 years of monthly payments tells us how much each person pays; we can then use that to determine the interest

OpenStudy (anonymous):

and theres two intrest rate 3% and 4%

OpenStudy (amistre64):

the interest rates might be used in some other approach, but lets do this our total cost is equal to , loan amount + interest paid 48M = 6500 + Mi 48 J = 6500 + Ji Mi - Ji is the difference between them

OpenStudy (amistre64):

(48M - 6500) -(48J - 6500) = difference in interest 48M - 6500 -48J + 6500 48M -48J 48(M -J) and we know the monthly payments for M and J

OpenStudy (anonymous):

so M and J are the monthy payments?

OpenStudy (amistre64):

yep

OpenStudy (anonymous):

so it would be 48(144-147)?

OpenStudy (amistre64):

yep

OpenStudy (anonymous):

it equals -144

OpenStudy (amistre64):

48*3 seems fair yes

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Thanks for the help did you want a medal

OpenStudy (amistre64):

medal is up to you ... i just wanted to help solve the problem

OpenStudy (amistre64):

good luck

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

was it correct?